Compounding annually means that there is one interest payment period per year.​ So, n​ = 5 and i​ = r​ = 0.08.

0J6g7WRD+GIF7wo5jkcVWXTox3F8UMZR46MI0fGk

UuHycdM5FNr7oeAFnVY+9artUVymLrnYIb4x67RF

jfb6bpyQMlenceSKwySRR4UCYGRMgRGyhAYKUNgp

​Therefore, rounded to the nearest cent​,

the final answer is _____________?
Answer #1

YES, INTEREST PAID ONLY ONCE IN A YEAR. FORMULA APPLIED IS ALSO CORRECT

FV = PV (1+r)n = 1000 (1+0.08)5 = 1000 X 1.4693280768 = 1469.3280768 = 1469.33

FINAL ANSWER : $1469.33